Assume that you are dealt four cards from a well-shuffled standard deck.

What is the probability of 4 aces given that you have at least 3 aces?
Answer #1

P(at least 3 aces)=P(3 ace and 1 other card)+P(4aces) =(4C3)*(48C1)/(52C4)+(4C4)*(48C0)/(52C4)

=4*48/270725+1*1/270725 =193/270725

therefore P(4 aces given at least 3 aces)=P(4 aces)/P( at least 3 aces)

=(1/270725)/(193/270725) =1/193